HUBBLE SPACE TELESCOPE DAILY REPORT #3068 PERIOD COVERED: 0000Z (UTC) 03/01/02 - 0000Z (UTC) 03/02/02 Daily Status Report as of 061/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.2 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.3 Completed STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.4 Completed Two Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.5 Completed Two Sets of STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.6 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.7 Completed Seven Sets of WF/PC-2 8942 (Cycle 10 Intflat Sweeps and Linearity Test) The WF/PC-2 was used to monitor the pixel-to-pixel flatfield response and provide a linearity check. These intflat sequences are done once during the year and the images will provide a backup database in the event of complete failure of the visflat lamp as well as allow monitoring of the gain ratios. There were no reported problems. 1.8 Completed STIS/CCD/MA1 9179 (Calibrating Convection Efficiency With Quasi-Molecular Features In Magnetic White Dwarfs) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe recently-discovered quasi-molecular absorption features of Lyman Alpha in UV spectra of magnetic white dwarfs taken with IUE. Hitherto, such features have only been known in non-magnetic objects where they offered a very accurate determination of the atmospheric parameters. There were no reported problems. 1.9 Completed Six Sets of WF/PC-2 8940 (Cycle 10 Earth Flats) The WF/PC-2 was used to monitor flatfield stability by obtaining sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed without incident. 1.10 Completed STISMA1/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.11 Completed STIS/CCD/MA2 9161 (The Ultraviolet Properties of Obscured QSOs) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to observe a subset of The Two Micron All-Sky Survey {2MASS}-discovered QSOs to: 1} determine if UV absorption- line features are present as seen in several other highly polarized AGN populations; 2} test relationships between X-ray and UV/optical attenuation, AGN spectral type, optical polarization, and detected UV absorption-line features; and 3} characterize the spectral energy distributions and UV properties of obscured QSOs. There were no problems. 1.12 Completed STIS/CCD/MA1/MA2 8919 (MAMA Sensitivity and Focus Monitor C10) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to monitor the sensitivity of each MAMA grating mode to detect any change due to contamination or other causes, and also to monitor the STIS focus in a spectroscopic and an imaging mode. All observations were successful. 1.13 Completed STIS/CCD 8907 (Spectroscopic Flats C10) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectrographic mode. There were no reported problems. 1.14 Completed WF/PC-2 9145 (A Snapshot Survey of the Optically Selected Type-2 Quasars) The WF/PC-2 was used to observe an identified population of emission-line objects in DPOSS, which can be plausibly interpreted as the long-sought type-2 quasars. They have high-ionization Seyfert-2 like spectra, but with narrow-line luminosities comparable to those of the luminous type-1 quasars in the same redshift range. This population may be a major contributor to the cosmic hard x-ray background. It is proposed to obtain multi-color images of a representative sample of these objects, in order to examine their morphology. We may be able to detect point-like nuclei which are not detectable in ground-based images, the dust lanes hiding them from our view, possible evidence for tidal interactions and the overall morphology of their hosts, etc. The proposal completed with no reported problems. 1.15 Completed HST 4779 (Vehicle Attitude And SI Configuration For Servicing Mission) In support of the Servicing Mission 3B, the vehicle must be at a +V3 sun pointing attitude. The science instruments must be in the launch configuration. This proposal defines the end of the last STScI SMS. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: Information Not Available Successful: Information Not Available Scheduled Re-acquisitions: Information Not Available Successful: Information Not Available 2.2 FHST Updates: Scheduled: Information Not Available Successful: Information Not Available 2.3 Operations Notes: STS-109 launched at 060/112201.985Z. Science operations terminated following the proposal described in 1.14. Currently the HST is in a Health and Safety load. The HST is scheduled to be captured by the Shuttle at approximately 062/2200Z. This is the last Daily Report until science operations resume following Servicing Mission 3B. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Resumption of normal science observations and calibrations following Servicing Mission 3B.. /CAW